# Chemical Cartography with APOGEE: Two-process Parameters and Residual Abundances for 288,789 Stars from Data Release 17

### v1

@@ 1 Introduction
As a first application of our expanded catalog, we examine the median residual abundances of 14 open clusters, nine globular clusters, and four dwarf satellites (the Large and Small Magellanic Clouds, the Sagittarius dwarf and stream, and Gaia Sausage/Enceladus). Taking advantage of medians for large samples allows us to robustly measure 0.05-dex level abundance differences between these stellar populations and MW field stars matched in overall CCSN and SNIa enrichment. The open cluster sample allows us to examine trends of individual elements with cluster age, similar to the analysis presented by Griffith et al. (2022) using GALAH data. The most important product of this paper is our catalog of two-process parameters and residual abundances, which can also be used to obtain [X/H] values corrected for systematics. We anticipate that this catalog will be useful for many investigations, including: characterizing additional enrichment sources such as AGB stars or rare supernovae; quantifying stochastic effects such as incomplete sampling of the initial mass function (IMF); revealing subtle differences in the abundance patterns of the bulge, radial and vertical zones of the disk, halo, and satellites; identifying groups with distinctive abundance patterns that could indicate a common birth environment; identifying trends with stellar age or binarity; defining chemically homogeneous samples for reconstructing the Galactic potential (Price-Whelan et al., 2021); uncovering trends with stellar age or binarity; and identifying rare outliers that could be signs of exotic physical processes or unusual measurement errors.
@@ 2 Data
[X/H] abundances are calculated from the sum of [X/Fe] (reported by either ASPCAP or BAWLAS) and [Fe/H] (reported by ASPCAP). For the uncertainty on [X/H], we use the reported [X/Fe] measurement uncertainty from ASPCAP (X_FE_ERR) or the empirical [X/Fe] uncertainty from BAWLAS (X_FE_ERR_EMP). ASPCAP estimates uncertainties empirically as a function of S/N, , and metallicity using repeated measurements of stars, and these uncertainties typically are larger than those reported by spectral model fitting (Jönsson et al., 2020, Section 5.4). BAWLAS reports two sets of uncertainties: a measurement uncertainty, determined from the scatter in line-by-line measurements of a given element, and and an empirical uncertainty, estimated from repeat observations (Hayes et al., 2022, Section 4.7). We use the BAWLAS empirical uncertainties because they are most similar in spirit to the ASPCAP uncertainties, though there are small differences in methodology (e.g., BAWLAS empirical uncertainties do not depend on SNR or overall metallicity, but do depend on individual abundances). Since each star uses the same [Fe/H] measurement in our calculation of [X/H], and we are primarily interested in the differential scatter between elements not between individual stars, we do not use the reported ASPCAP [Fe/H] uncertainty except when considering . For the combined element C+N, we follow W22 and take the [C/Fe] uncertainty as the [(C+N)/H] uncertainty.
@@ 4.2 Two-process Amplitudes, and
In Figure 8, we plot the distribution of best-fit vs. values. The high-Ia (in blue) and low-Ia (in red) populations separate, similar to a typical [Mg/Fe] vs. [Fe/H] plot but more closely analogous to [Fe/Mg] vs. [Mg/H], with providing a measure of metallicity as traced by CCSN elements. The division of the two populations at reflects the definition of the populations (Equation 2). Unlike the flat plateau at for the low-Ia population in an [Fe/Mg] vs. [Mg/H] plot, the low-Ia population shows a steady rise in as increases. The high-Ia population shows a similar but steeper increase in before plateauing at . The rise in both populations traces an increase in time-delayed SNIa enrichment over time, and the shallower rise in the high-Ia trend plausibly results from higher star formation efficiency, which allows the population to reach higher within the characteristic SNIa delay time Gyr. The flat plateau at suggests that these stars form when the CCSN and SNIa rates are approximately equal to each other, producing near-solar abundance ratios. These general trends are more obvious in the calibration sample, which consists only of MW disk stars. The high feature at is not seen in the disk and consists primarily of dwarf galaxy members (Section 5.3). In detail, the appearance of this diagram depends on our key model assumption that [Mg/Fe] = 0.3 is the ratio produced by pure CCSN enrichment () (Section 4.1).
@@ 5.1 Open Clusters
Ce enhancement sometimes arises in concert with C+N enhancement as a consequence of binary mass transfer (e.g., Smith et al. 2000, and discussed in the context of residual abundances in W22). However, we expect this phenomenon to affect only a small fraction of stars, so the open cluster Ce enhancement is likely present in birth abundances. Outside clusters, Casali et al. (2023) studied low- stars with astroseismic ages and found that younger stars have higher [Ce/Fe] at fixed [Fe/H] and higher [Ce/] at fixed radial position. In residual abundances, W22 find a trend of enhanced Ce in young APOGEE field stars, and Griffith et al. (2022) find a similar trend for Y and Ba in GALAH, so it seems probable that both cluster and field stars reflect a general increase of AGB enrichment in the ISM at late times, relative to SNIa enrichment. The C+N enhancement in our two youngest clusters might be a consequence of such extra AGB enrichment, and it is tempting to ascribe the elevated levels of Na, S, and Cu in these clusters to the same source. However, the theoretically predicted AGB contribution to these three elements is small, in contrast to Ce and C+N (see, e.g., Fig. 10 of Andrews et al. 2017 and Fig. 13 of Rybizki et al. 2017).
@@
Na overabundance in evolved stars can be caused by mixing during the first dredge-up in low- to intermediate-mass red giants (Smiljanic et al., 2016, and references therein). Our observed Na enhancement in only the youngest clusters could reflect this phenomenon since their evolved stars (which make up our sample) are more massive compared to older clusters. In contrast to Griffith et al. (2022), we do not see any notable Ca or K enhancement in any clusters; they also did not see any enhancement in Na. We also observe mild Cu enhancement in our youngest clusters where they observed depletion. The other elements showing enhancement (S and C+N) were not considered in Griffith et al. (2022). An advantage of APOGEE for investigating young clusters is that the abundance determinations for red giants are less likely to be affected by rotational line broadening than determinations from main sequence stars that dominate GALAH cluster samples.
@@
ESO 211-03 and NGC 2158 show significant (0.2 dex) depletion in Co. To investigate this depletion, we compare the strongest Co line in the observed APOGEE spectra to synthetic spectra. Using the 1D spectral synthesis code Korg (Wheeler et al., 2023a, b) with the APOGEE DR17 line list (most similar to that of the DR16 line list, described in Smith et al. 2021), we generate two synthetic spectra for every star in each cluster: one spectrum with reported ASPCAP abundances and stellar parameters, and one spectrum where only the Co abundance is changed to that predicted by the two-process model. We note that we use the uncalibrated, best-fit ASPCAP spectroscopic and as Korg inputs to synthesize the spectra. The synthetic spectra are downsampled to APOGEE resolution, then normalized by dividing out a moving mean, where at each pixel the mean is calculated using a bandwidth of 25Å. For the observed spectra, we use the combined, but not normalized, APOGEE spectra in the apStar/asStar files and normalize using the same moving mean procedure with a weighted mean to account for flux uncertainties. Finally, we calculate the pixel-by-pixel median of the normalized spectra over every star in the cluster. As illustrated in Figure 14, we find that the median observed spectrum is a closer match to the median synthetic spectrum with reported ASPCAP Co abundances in both ESO 211-03 and NGC 2158. This indicates that the Co prediction of the two-process model results in a poorer fit, and thus that the observed depletion is most likely real. However, the physical origin of this Co depletion is unclear. An alternative explanation for the significant mismatch between the observed spectra and two-process prediction is that the spectra are affected by small errors in sky subtraction: the 16762Å Co line falls on a telluric feature in both ESO 211-03 and NGC 2158. In either case, we show that our median residual abundance analysis can reveal real differences in spectra of stellar populations. A caveat of this analysis is that while we use identical line lists and model atmospheres, Korg is not the same spectral synthesis code used to generate the ASPCAP spectral grids for DR17, so we are in effect generating new spectra to compare to APOGEE.
@@
NGC 1245, NGC 7789, and NGC 2420 show a 0.1-0.2 dex enhancement in median V. We performed a similar analysis to that of the Co depletion for ASPCAP V (chosen over BAWLAS due to a higher number of measurements) and found that the median observed APOGEE spectra more closely matched that of the median synthetic spectrum with predicted V abundances for all three clusters. From this result we conclude that the V enhancements seen in the panels of Figure 11 are likely not real, as the median line depth is consistent with the two-process prediction (i.e., the residual should be close to 0). The V residuals show some of the largest variations between individual stars in the majority of the clusters, so there are likely large uncertainties in the measurements as well.
@@
Finally, we check the enhancements for the remaining single elements showing age trends, S, Ce, and Cu, using the median spectrum analysis in Korg. For Ce, we performed this analysis on five clusters showing 0.1 dex median enhancement (NGC 6705, NGC 1245, ESO 211-03, NGC 2420, and NGC 2204) and used the ASPCAP abundances to increase the number of measurements in each cluster. The median spectra show clear enhancements in Ce, similarly to the Co case (Figure 14) but to a smaller degree, in all clusters except NGC 6705. NGC 6705 suffered similar continuum matching issues as in the Na case. In S, we found that the median observed spectra were actually shallower than even the two-process prediction i.e., the clusters have less S than predicted. Upon further investigation, it is likely that several line blends in S in combination with the initial fitting of S with the other alpha elements resulted in ASPCAP overestimating the S abundance. However, in NGC 6705 and Berkeley 85, using the smaller number of BAWLAS measurements of S, whose methods are intended to better measure abundances of elements suffering from blended lines, yields very similar results to using ASPCAP. For Cu, we examined NGC 6705 and Berkeley 85, and we find a similar result to S, where the median observed spectra were shallower than the median synthetic spectra. However, the 0.1 dex median Cu enhancements seen in these clusters is insufficient to generate a clear difference between the median synthetic spectra generated with measured or predicted (two-process) Cu at APOGEE resolution, so it would have been difficult to determine which is closer to the data regardless.
@@ 5.2.3 Iron-peak Elements
The two Fe-peak elements (Fe and Ni) used to fit the two-process amplitudes generally have a very small, if any, median residual in our GC sample except for Cen, which shows a 0.2 dex depletion in Ni. Cen also has 0.1-0.3 dex depletions (depending on the cluster/element) in the other Fe-peak elements V, Mn, and Co in addition to the aforementioned Ni; this chemical pattern appears more similar to our sample of massive dwarf satellite galaxies (Figure 18) than the other GCs so we discuss its potential origins in Section 5.3.4. M4 has an 0.2 dex median residual in Cr, while NGC 6388 and NGC 6380 show smaller 0.1 dex Cr residuals. NGC 6388 and NGC 6380 also have 0.1-0.2 dex enhancements in Co. NGC 6388, NGC 6380, and NGC 6760 show 0.1-0.2 dex enhancement in V. V also shows large star-by-star scatter in its residuals; however, because the star-by-star deviations are present and large in other stellar populations (e.g. Figures 11, 12, 18), this is unlikely to be a signature of multiple populations, unlike in the case of Na. Furthermore, M107 shows a V enhancement based on the BAWLAS measurement and a V depletion based on the ASPCAP measurements, which suggests that the variation in V is likely due to large measurement errors. This effect is present but smaller in Cen, where the median BAWLAS residual is just barely positive while the median ASPCAP residual shows a small depletion of 0.1 dex. Only four GCs in our sample (47 Tuc, M4, M71, and Cen) have any Cu measurements. Of these four GCs, only 47 Tuc has more than six measurements, and it has [Cu/H]0.
@@ 5.3 Milky Way Satellite Galaxies
We plot [Fe/Mg] vs. [Mg/H] for the member stars in each of the satellite galaxies in our sample in the left panels of Figure 18. All four galaxies lie on or above the median high-Ia sequence, indicating extended star formation histories that allowed the most metal-rich stars in these systems to be enriched by many generations of delayed SNIa contributions. GSE shows the greatest number of stars below the high-Ia median sequence, which makes sense given that its star formation was most likely cut off 10 Gyr ago during its infall into the MW (e.g., Helmi et al., 2018; Gallart et al., 2019; Vincenzo et al., 2019). A “wall” of stars exists at the leftmost edge of the panels, right at the sample metallicity cutoff. When compared to the [Fe/Mg] vs. [Mg/H] plots in Hasselquist et al. (2021) (Figures 6 and 7), it is clear that our work samples only the most metal-rich end of these satellite galaxies’ metallicity distributions.
@@ 5.3.4 Common Trends
The residual abundance diagram (Figure 18) provides a compact way to compare and contrast multi-dimensional enrichment patterns among satellites and between satellites and the disk, controlling for differences associated with the overall metallicity and SNIa/CCSN ratio, and leveraging the ability of median statistics to reveal systematic differences below the level of individual star measurement errors. Overall, we find that the four massive MW satellite galaxies represented in our data (LMC, SMC, Sgr, and GSE) show surprisingly similar residual abundance patterns despite their different star formation environments and enrichment histories, as well as only sampling from the most metal-rich stars in these systems due to our sample selection constraints. Combining the KPM methodology of Griffith et al. (2023) with the calibration strategy developed here would allow these comparisons to be extended to the lower metallicity populations of dwarf satellites, tidal streams, and the in-situ halo.

### v2

@@ 1 Introduction
As a first application of our expanded catalog, we examine the median residual abundances of 14 open clusters, nine globular clusters, and four dwarf satellites (the Large and Small Magellanic Clouds, the Sagittarius dwarf and stream, and Gaia Sausage/Enceladus). Taking advantage of medians for large samples allows us to robustly measure 0.05-dex level abundance differences between these stellar populations and MW field stars matched in overall CCSN and SNIa enrichment. The open cluster sample allows us to examine trends of individual elements with cluster age, similar to the analysis presented by Griffith et al. (2022) using GALAH data. The most important product of this paper is our catalog of two-process parameters and residual abundances, which can also be used to obtain [X/H] values corrected for systematics. We anticipate that this catalog will be useful for many investigations, including: characterizing additional enrichment sources such as AGB stars or rare supernovae; quantifying stochastic effects such as incomplete sampling of the initial mass function (IMF); revealing subtle differences in the abundance patterns of the bulge, radial and vertical zones of the disk, halo, and satellites; identifying groups with distinctive abundance patterns that could indicate a common birth environment; uncovering trends with stellar age or binarity; defining chemically homogeneous samples for reconstructing the Galactic potential (Price-Whelan et al., 2021); and identifying rare outliers that could be signs of exotic physical processes or unusual measurement errors.
@@ 2 Data
[X/H] abundances are calculated from the sum of [X/Fe] (reported by either ASPCAP or BAWLAS) and [Fe/H] (reported by ASPCAP). For the uncertainty on [X/H], we use the reported [X/Fe] measurement uncertainty from ASPCAP (X_FE_ERR) or the empirical [X/Fe] uncertainty from BAWLAS (X_FE_ERR_EMP). We are primarily interested in the differential scatter between elements not between individual stars, and each star uses the same [Fe/H] measurement in our calculation of [X/H], so although [Fe/H] has its own reported ASPCAP uncertainty, the impact of the [Fe/H] uncertainty is completely correlated with [X/H], and we therefore do not consider it in our [X/H] uncertainties except in the case . For the combined element C+N, we follow W22 and take the [C/Fe] uncertainty as the [(C+N)/H] uncertainty. ASPCAP estimates uncertainties empirically as a function of S/N, , and metallicity using repeated measurements of stars, and these uncertainties typically are larger than those reported by spectral model fitting (Jönsson et al., 2020, Section 5.4). BAWLAS reports two sets of uncertainties: a measurement uncertainty, determined from the scatter in line-by-line measurements of a given element, and and an empirical uncertainty, estimated from repeat observations (Hayes et al., 2022, Section 4.7). We use the BAWLAS empirical uncertainties because they are most similar in spirit to the ASPCAP uncertainties, though there are small differences in methodology (e.g., BAWLAS empirical uncertainties do not depend on SNR or overall metallicity, but do depend on individual abundances).
@@ 4.2 Two-process Amplitudes, and
In Figure 8, we plot the distribution of best-fit vs. values. The high-Ia (in blue) and low-Ia (in red) populations separate, similar to a typical [Mg/Fe] vs. [Fe/H] plot but more closely analogous to [Fe/Mg] vs. [Mg/H], with providing a measure of metallicity as traced by CCSN elements. The division of the two populations at reflects the definition of the populations (Equation 2). Unlike the flat plateau at for the low-Ia population in an [Fe/Mg] vs. [Mg/H] plot, the low-Ia population shows a steady rise in as increases. The high-Ia population shows a similar but steeper increase in before plateauing at . The rise in both populations traces an increase in time-delayed SNIa enrichment over time, and the shallower rise in the high-Ia trend plausibly results from higher star formation efficiency, which allows the population to reach higher within the characteristic SNIa delay time Gyr. The flat plateau at suggests that these stars form when the CCSN and SNIa rates are approximately equal to each other, producing near-solar abundance ratios. These general trends are more obvious in the calibration sample, which consists only of MW disk stars. The high feature at is not seen in the disk and consists primarily of dwarf galaxy members (Section 5.3). The detailed appearance of this diagram depends on our key model assumption that [Mg/Fe] = 0.3 is the ratio produced by pure CCSN enrichment () (Section 4.1).
@@ 5.1 Open Clusters
Ce enhancement sometimes arises in concert with C+N enhancement as a consequence of binary mass transfer (e.g., Smith et al. 2000, and discussed in the context of residual abundances in W22). However, we expect this phenomenon to affect only a small fraction of stars, so the open cluster Ce enhancement is likely present in birth abundances. Outside clusters, Casali et al. (2023) studied low- stars with astroseismic ages and found that younger stars have higher [Ce/Fe] at fixed [Fe/H] and higher [Ce/] at fixed radial position. In residual abundances, W22 find a trend of enhanced Ce in young APOGEE field stars, and Griffith et al. (2022) find a similar trend for Y and Ba in GALAH, so a plausible interpretation is that both cluster and field stars reflect a general increase of AGB enrichment in the ISM at late times, relative to SNIa enrichment. The C+N enhancement in our two youngest clusters might be a consequence of such extra AGB enrichment, and it is tempting to ascribe the elevated levels of Na, S, and Cu in these clusters to the same source. However, the theoretically predicted AGB contribution to these three elements is small, in contrast to Ce and C+N (see, e.g., Fig. 10 of Andrews et al. 2017 and Fig. 13 of Rybizki et al. 2017).
@@
Na overabundance in evolved stars can be caused by mixing during the first dredge-up in low- to intermediate-mass red giants (Smiljanic et al., 2016, and references therein). Our observed Na enhancement in only the youngest clusters could reflect this phenomenon since their evolved stars (which make up our sample) are more massive compared to older clusters. In contrast to Griffith et al. (2022), we do not see any notable Ca or K enhancement in any clusters; however, these enhancements are strongest in clusters younger than 300 Myr, and we do not have any clusters of this age in our sample. Conversely, Griffith et al. (2022) did not see any enhancement in Na. We also observe mild Cu enhancement in our youngest clusters where they observed depletion. The other elements showing enhancement (S and C+N) were not considered in Griffith et al. (2022). An advantage of APOGEE for investigating young clusters is that the abundance determinations for red giants are less likely to be affected by rotational line broadening than determinations from main sequence stars that dominate GALAH cluster samples.
@@
ESO 211-03 and NGC 2158 show significant (0.2 dex) depletion in Co. To investigate this depletion, we compare the strongest Co line in the observed APOGEE spectra to synthetic spectra. Using the 1D spectral synthesis code Korg (Wheeler et al., 2023a, b) with the APOGEE DR17 line list (most similar to that of the DR16 line list, described in Smith et al. 2021), we generate two synthetic spectra for every star in each cluster: one spectrum with reported ASPCAP abundances and stellar parameters, and one spectrum where only the Co abundance is changed to that predicted by the two-process model. We note that we use the uncalibrated, best-fit ASPCAP spectroscopic and as Korg inputs to synthesize the spectra. The synthetic spectra are downsampled to APOGEE resolution, then normalized by dividing out a moving mean, where at each pixel the mean is calculated using a bandwidth of 25Å. For the observed spectra, we use the combined, but not normalized, APOGEE spectra in the apStar/asStar files and normalize using the same moving mean procedure with a weighted mean to account for flux uncertainties. Finally, we calculate the pixel-by-pixel median of the normalized spectra over every star in the cluster. As illustrated in Figure 14, we find that the median observed spectrum is a closer match to the median synthetic spectrum with reported ASPCAP Co abundances in both ESO 211-03 and NGC 2158. This indicates that the Co prediction from the two-process model results in a poorer fit, and thus that the observed depletion is most likely real. However, the physical origin of this Co depletion is unclear. An alternative explanation for the significant mismatch between the observed spectra and two-process prediction is that the spectra are affected by small errors in sky subtraction: the 16762Å Co line falls on a telluric feature in both ESO 211-03 and NGC 2158. In either case, we show that our median residual abundance analysis can reveal real differences in spectra of stellar populations. A caveat of this analysis is that while we use identical line lists and model atmospheres, Korg is not the same spectral synthesis code used to generate the ASPCAP spectral grids for DR17, so we are in effect generating new spectra to compare to APOGEE.
@@
NGC 1245, NGC 7789, and NGC 2420 show a 0.1-0.2 dex enhancement in median V in Figure 11. We performed a similar analysis to that of the Co depletion for ASPCAP V (chosen over BAWLAS due to a higher number of measurements) and found that the median observed APOGEE spectra more closely matched that of the median synthetic spectrum with predicted two-process V abundances for all three clusters. From this result we conclude that the V enhancements seen in Figure 11 are likely not real because the median line depth of the observed spectra is consistent with the two-process prediction (i.e., the residual should be close to 0). The V residuals show some of the largest variations between individual stars in the majority of the clusters, so there are likely large uncertainties in the measurements as well.
@@
Finally, we check the enhancements for the remaining single elements showing age trends, S, Ce, and Cu, using the median spectrum analysis in Korg. For Ce, we performed this analysis on five clusters showing 0.1 dex median enhancement (NGC 6705, NGC 1245, ESO 211-03, NGC 2420, and NGC 2204) and used the ASPCAP abundances to increase the number of measurements in each cluster. The median spectra show clear enhancements in Ce, similarly to the Co case (Figure 14) but to a smaller degree, in all clusters except NGC 6705. NGC 6705 suffered similar continuum matching issues as in the Na case. In S, we found that the median observed spectra were actually shallower than even the two-process prediction i.e., the clusters have less S than predicted. Upon further investigation, it is likely that several line blends in S in combination with the initial fitting of S with the other elements resulted in ASPCAP overestimating the S abundance. However, in NGC 6705 and Berkeley 85, using the smaller number of BAWLAS measurements of S, whose methods are intended to better measure abundances of elements suffering from blended lines, yields very similar results to using ASPCAP. For Cu, we examined NGC 6705 and Berkeley 85, and we find a similar result to S, where the median observed spectra were shallower than the median synthetic spectra. However, the 0.1 dex median Cu enhancements seen in these clusters is insufficient to generate a clear difference between the median synthetic spectra generated with measured or predicted (two-process) Cu at APOGEE resolution, so it would have been difficult to determine which is closer to the data regardless.
@@ 5.2.3 Iron-peak Elements
The two Fe-peak elements (Fe and Ni) used to fit the two-process amplitudes generally have a very small, if any, median residual in our GC sample except for Cen, which shows a 0.2 dex depletion in Ni. Cen also has 0.1-0.2 dex depletions in the other Fe-peak elements V, Mn, and Co in addition to the aforementioned Ni; this chemical pattern in the Fe-peak elements appears more similar to our sample of massive dwarf satellite galaxies (Figure 18) than the other GCs so we discuss its potential origins in Section 5.3.4. M4 has an 0.2 dex median residual in Cr, while NGC 6388 and NGC 6380 show smaller 0.1 dex Cr residuals. NGC 6388 and NGC 6380 also have 0.1-0.2 dex enhancements in Co. NGC 6388, NGC 6380, and NGC 6760 show 0.1-0.2 dex enhancement in V. V also shows large star-by-star scatter in its residuals; however, because the star-by-star deviations are present and large in other stellar populations (e.g. Figures 11, 12, 18), this is unlikely to be a signature of multiple populations, unlike in the case of Na. Furthermore, M107 shows a V enhancement based on the BAWLAS measurement and a V depletion based on the ASPCAP measurements, which suggests that the variation in V is likely due to large measurement errors. This effect is present but smaller in Cen, where the median BAWLAS residual is just barely positive while the median ASPCAP residual shows a small depletion of 0.1 dex. Only four GCs in our sample (47 Tuc, M4, M71, and Cen) have any Cu measurements. Of these four GCs, only 47 Tuc has more than six measurements, and it has [Cu/H]0.
@@ 5.3 Milky Way Satellite Galaxies
We plot [Fe/Mg] vs. [Mg/H] for the member stars in each of the satellite galaxies in our sample in the left panels of Figure 18. All four galaxies lie on or above the median high-Ia sequence, indicating extended star formation histories: the gas that formed the most metal-rich stars in these systems had time to be enriched by many generations of delayed SNIa, thus increasing the relative amount of Fe to Mg. GSE shows the greatest number of stars below the high-Ia median sequence, which makes sense given that its star formation was most likely cut off 10 Gyr ago during its infall into the MW (e.g., Helmi et al., 2018; Gallart et al., 2019; Vincenzo et al., 2019), so later SNIa contributions were not incorporated into the most metal-rich stars. A “wall” of stars exists at the leftmost edge of the panels, right at the sample metallicity cutoff. The metallicity range we can probe in this study is limited by the metallicity range of the MW disk calibration sample, chosen to match W22 and such that we observe clear [/Fe] bimodality over its metallicity range. When compared to the [Fe/Mg] vs. [Mg/H] plots in Hasselquist et al. (2021) (Figures 6 and 7), it is clear that our work samples only the most metal-rich end of these satellite galaxies’ metallicity distributions.
@@ 5.3.4 Common Trends
The residual abundance diagram (Figure 18) provides a compact way to compare and contrast multi-dimensional enrichment patterns among satellites and between satellites and the disk, controlling for differences associated with the overall metallicity and SNIa/CCSN ratio, and leveraging the ability of median statistics to reveal systematic differences below the level of individual star measurement errors. Overall, we find that the four massive MW satellite galaxies represented in our data (LMC, SMC, Sgr, and GSE) show surprisingly similar residual abundance patterns despite their different star formation environments and enrichment histories, as well as only sampling from the most metal-rich stars in these systems due to our sample selection constraints. Combining the -process model (KPM) methodology of Griffith et al. (2023) with the calibration strategy developed here would allow these comparisons to be extended to the lower metallicity populations of dwarf satellites, tidal streams, and the in-situ halo.